# Reducing Spatial Discretization Error on Coarse CFD Simulations Using an OpenFOAM-Embedded Deep Learning Framework

### v2

@@ Abstract
We propose a method for reducing the spatial discretization error of coarse computational fluid dynamics (CFD) problems by enhancing the quality of low-resolution simulations using a deep learning model fed with high-quality data. We substitute the default differencing scheme for the convection term by a feed-forward neural network that interpolates velocities from cell centers to face values to produce velocities that approximate the fine-mesh data well. The deep learning framework incorporates the open-source CFD code OpenFOAM, resulting in an end-to-end differentiable model. We automatically differentiate the CFD physics using a discrete adjoint code version. We present a fast communication method between TensorFlow (Python) and OpenFOAM (c++) that accelerates the training process. We applied the model to the flow past a square cylinder problem, reducing the error to about 50% for simulations outside the training distribution compared to the traditional solver in the $x$- and $y$-velocity components using an 8x coarser mesh. The training is affordable in terms of time and data samples since the architecture exploits the local features of the physics while generating stable predictions for mid-term simulations.
@@ 1 Introduction
Over the last decade, machine learning (ML) has arisen as a crucial technology for enhancing the solving of partial differential equations (PDEs) [1]. The field of fluid dynamics has been part of this revolution. Computational fluid dynamics (CFD) witnessed the introduction of novel learning techniques [2]. Non-linear PDEs describe fluid motion, usually showing chaotic spatiotemporal dependencies and structures with a wide range of scales, making direct numerical simulation (DNS) unfeasible in most cases. Alternatives like Large Eddy Simulations (LES) [3, 4] or Reynolds-Averaged Navier-Stokes (RANS) [5] exist, where only large structures are solved while the rest are modeled. However, the resolution process can be complex and time-consuming, even with LES or RANS.
@@
Enhancing CFD simulations through ML can accelerate the computations without compromising accuracy or produce more accurate results within the same computational time. For this purpose, the ML model can completely substitute the traditional CFD solver, leading to reduced-order models (ROMs) or partially enhancing some specific components’ performance. The ROMs exhibit accelerated time-to-solution but pay the cost of low fidelity and a narrow range of applicability [6, 7, 8, 9]. ROMs use linear data compression techniques like dynamic-mode decomposition (DMD) [10] and principal-component analysis (PCA) [11], or non-linear compression like autoencoders [12]. When substituting a component of the traditional solver, the typical approach consists of replacing the turbulent closure model  [13], which is one of the significant sources of error, with an ML model. Thus, Ling et al. [14] outperform the traditional solvers using a deep neural network for RANS turbulence modeling. Alternatively, Obiols-Sales et al. [15] speed up the convergence of RANS simulations by combining iterations of a convolutional neural network (CNN) and the traditional solver. Ajuria et al. [16] accelerate the algorithm by substituting the Jacobi solver with a CNN for the Poisson equation, one of the most expensive parts of the traditional numerical solver.
@@
Ideally, the ML model should enhance the calculation performance within multiple initial/boundary conditions, be accurate for unseen long-term CFD simulations, and use an acceptable amount of data and time for training. For example, Kim et al. [17] improved the precision of the Finite Volume Method (FVM) [18] combining predictions of a pre-trained neural network [19] and of a traditional solver. Finally, another promising approach consists on reducing the discretization error on coarse meshes [20], applying a method called learned discretizations [21]; for example, [22] estimates spatiotemporal operators on low-resolution grids through CNNs for accelerating DNS and LES simulations, learning from fine-resolution data. This technique calls the CFD solver during training in a solver-in-the-loop [23] ML model. This coupling demands an end-to-end differentiable framework; thus, translating the fluid solver to an automatic differentiable language is a stringent requirement [24]. This limitation reduces the applicability of this kind of method to simple cases. In addition, using CNN implies a considerably large dataset, making its training process demanding.
@@
We extend the learned discretization technique to a real case with non-trivial geometries and more complex flow dynamics, such as the flow past a square cylinder problem [25]. In these convection-dominated flows, the discretization of the convective term is a critical source of error where the differencing scheme relating the cell´s face and center values plays a major role in stability and precision. Numerous scheme versions may be found in the literature depending on the conditions of the simulated case. We reduce the spatial discretization error on coarse meshes for unsteady RANS [5] simulations by replacing the traditional differencing scheme with a feed-forward neural network, which learns from high-resolution data.
@@
In contrast to previous works, we embed the open-source CFD code OpenFOAM [26] in the training loop to exploit the vast range of physics and numerical solvers this program has implemented. Besides, our neural network’s architecture exploits the local features of the flow to speed up the training process. We also develop a fast communication method between the ML framework and the CFD code to accelerate the training process. Our algorithm applies the technique to our case study, but it is easily generalisable to any fluid case simulated in OpenFOAM.
@@
Our algorithm imitates a machine learning technique commonly applied in the image processing field called super-resolution [27]. The method increases the resolution of low-quality images based on high-quality data employing data-driven models. Our model enforces physical constraints of the governing equations while enriching certain parts of the traditional solver with high-resolution data. Its main limitation resides in the necessity of data with an image format, that is, structured data generated from Cartesian meshes with cells of the same size.
@@ 2 Problem description
The flow past a square cylinder is a well-known engineering problem with multiple industrial applications in offshore platforms [28], bridge piers [29], high-rise buildings [30], or coastal engineering [31]. The state-of-the-art of this problem is well-developed with numerous numerical [32, 33, 34] and experimental studies [35, 36, 37], leading to a case that balances industry and research. There is no general convention between authors on the definition of the computational domain and boundary conditions, so we define our 2D problem as described in Figure 1 following the recommendations in [38], where $D$ is the side of the square column.
@@
According to the non-dimensional Reynolds number, which measures the ratio between inertial and viscous forces, we can identify different flow regimes [38, 32]. Most applications are interested in transition to turbulence or fully turbulent regimes, where attached vortex shedding appears in the wake of the obstacle and forms the von Karman street vortex effect. These streets can be more or less stable depending on the turbulence grade of the flow, resulting in unsteady numerical simulations.
We fix the diameter of the square column $D=1$ m, an inlet velocity of 1 m/s, and a kinematic viscosity $\nu=2\cdot 10^{-6}$ m/s2, obtaining a Reynolds number $Re=5\cdot 10^{5}$. Hence, the flow dynamic is completely turbulent, making it necessary to use a mesh that is fine enough to describe the physics with fidelity. We define our fine spatial discretization $f$ as a structured mesh formed by 303,146 quadrilateral cells of the same size. We establish a fixed temporal step of 0.005 s, obtaining a Courant-Friedrichs-Lewy (CFL) number [39] equals 0.32. This mesh generates the training data since it recovers adequately the von Karman effect, although at an increased computational cost. Upsampling this mesh isotropically, we define our coarse discretization $c$, which produces low-accuracy results. We apply our method to this mesh to enhance its solution from a learned model fed by high-accurate data. Figure 2 depicts a comparison between both discretizations.
Figure 2: Comparison of the same data in two resolutions. (Left) High-resolution data from the fine mesh. (Right) Low-resolution data from the x16 coarser mesh.
@@
From now on, we will present different operators and variables with the subscripts $c$ (coarse) or $f$ (fine), indicating to which spatial discretization we apply them. For the transformation of fine data into coarse dimensions, we introduce the projection operator $\mathcal{P}_{f}^{c}:\mathbb{R}^{m_{f}}\rightarrow\mathbb{R}^{m_{c}}$, which is applied to a generic scalar-valued field $\phi$ leads to:
$\phi_{c,f}=\mathcal{P}_{f}^{c}(\phi_{f})=\frac{1}{N}\sum_{j=1}^{N}\phi_{f}^{i}(j)\qquad\text{with $i\in\{1,...,m_{c}\}$},$ (1)
@@
where the subscript $c,f$ denotes the fine data projected onto the coarse dimension, $N$ is the number of fine cells inside a coarse cell $i$, and $m_{c}$ is the number of cells of the destination coarse mesh. This operator runs through each cell of the coarse mesh and computes the arithmetic mean of the fine data contained within the considered cell. With this conservative averaging, we ensure that the resulting projected field fulfills continuity.
@@ 3.1 Governing equations
The system of PDEs describing the dynamics of an incompressible fluid flow is called the Navier-Stokes equations (NSE). We are solving a Reynolds-averaged version of these equations, where all turbulence scales are modeled. Thus, the instantaneous values of the involved fields are decomposed into a mean component and a fluctuating one:
@@
3.2 Resolution algorithm
The complex PDE system describing the problem’s physics requires an ad-hoc resolution process. Multiple optimized algorithms exist depending on the fluid flow conditions; we use an operator-splitting method called Pressure IMplicit for Pressure-Linked Equations (PIMPLE [18]) for this case. This method is a predictor-corrector algorithm, formed by four main stages as described in Figure 3: (1) the velocity predictor, (2) the pressure equation stage, (3) the velocity corrector, and (4) the turbulence equations stage.
@@ 3.2 Resolution algorithm
First, we construct a linearized momentum equation derived from (4) using the FVM and taking as input the variables at time $t$. The solution of this equation is the predicted mean velocity $\bar{\textbf{u}}^{*}$, which does not satisfy continuity (3). During the second step, we form and solve a Poisson equation by combining (3)-(4), leading to a predicted mean pressure $\bar{p}^{*}$. During stage (3), the algorithm corrects the mean velocity based on the expected pressure computed in the previous step. This stage is repeated in the correction loop until the corrected mean velocity $\bar{\textbf{u}}^{**}$ is conservative. At the end of this stage, we obtain a candidate mean velocity and a candidate mean pressure, which can be sent back to stage (1) in case the convergence criteria are not satisfied. In that case, we would be restarting the PIMPLE loop taking $\bar{\textbf{u}}^{**}$ and $\bar{p}^{*}$ as the new input of the iteration. Once convergence is reached, we obtain the mean velocity and mean pressure for the time $t+1$. Finally, we jump into the final stage, where we solve the turbulence equations (5)-(6), producing $k$ and $\omega$ at time $t+1$.
@@ 3.3 Differencing schemes
As explained previously, the first step of the solver consists of linearizing the momentum equation to obtain a system of linear equations, which can be generalized for a cell $p$ as:
@@
where $N$ is the selected stencil around face $F$ and $w$ are the interpolation weights. A differencing scheme is an algebraic expression for relating a cell’s face and center values based on some constant interpolation weights. Experts have developed numerous differencing schemes (e.g., Upwind, QUICK, TVD schemes [18]) depending on the flow condition, resulting in variations of the stencil size and the interpolation weights. However, no general approach has been reached. For this reason, its choice is case-dependent, producing different results in accuracy and stability. Finally, the mean face velocity is not computed explicitly; in contrast, the interpolation weights $w_{i}$ contribute to the matrix of coefficients through $a_{p}$ and $a_{nb}$ of (7), following an implicit approach.
@@ 4.1 Formal description
Our problem is completely turbulent, so the convective contribution is the most crucial term of (4). The use of an appropriate mesh is essential for discretizing this term since a coarser mesh would considerably impact the accuracy due to its spatial discretization error. In particular, when we discretize the convective term using FVM, we obtain the following expression:
@@
As explained, the critical interpolation is carried out by a differencing scheme. Our approach consists of substituting the differencing scheme used for the convective term with a custom version based on data-driven discretizations [21]. Thus, we replace the constant interpolation weights $w_{i}$ of (8) by a vector of optimized interpolation weights $\widetilde{\textbf{w}}_{i}$ generated by a feed-forward neural network. We can therefore define a modified inverse operator ~$\widetilde{\mathcal{I}}:\mathbb{R}^{2N}\rightarrow\mathbb{R}^{2}$, such that:
@@
where $\theta$ are the parameters of the neural network. In contrast to traditional differencing schemes, our approach takes the mean velocity and mean pressure as inputs and returns a vector of interpolation weights that are space and time-dependent. We should emphasize both $\mathcal{I}$ and $\widetilde{\mathcal{I}}$ are local operators, which means they compute the value at face $F$ based only on its neighborhood. However, we need to apply the local operators to all the internal faces of the domain, resulting in the mean face velocities of the domain $\bar{\textbf{u}}_{F(n)}$ as:
@@
$\textbf{s}^{t}:=\bigcup^{m}_{j=1}\{\bar{\textbf{u}}^{t}_{(j)},\bar{p}^{t}_{(j)},k^{t}_{(j)},\omega^{t}_{(j)})\},$ (13)
being $m$ the number of cells of the mesh. We fully define our model by composing these operators ($\mathcal{F}\circ\mathcal{I})(\textbf{s}^{t})$ and applying them to a spatial discretization. In particular, we run the learned model on coarse meshes, which we indicate by the subscript $c$. We train it in such a way that the generated interpolation weights produce a mean velocity $\bar{\textbf{u}}_{c}$ as close as possible to the mean velocity from the fine discretization projected onto the coarse one $\bar{\textbf{u}}_{c,f}$. Thus, we can formulate the training as a depreciation of the following problem:
@@
Our method, called Deep-Convection (DC), tries to mimic the super-resolution techniques employed in image processing, which augment the resolution of low-quality images based on high-quality data [27]. With this approach, the primary benefit is that it maintains the physical constraints of the governing equations while exploiting the benefits of a data-driven model for enriching certain parts of the traditional solver with high-resolution data.
@@
where $\phi_{c}$ is a general scalar-valued field computed using the coarse discretization and $\phi_{c,f}$ is the projection of the fine data $\phi_{f}$ onto the coarse mesh $c$. In this case, $\phi_{c,f}$ is the training data or the ground truth. Due to its multi-step feature, we gather the solver prediction results over several accumulated timesteps $T$ and then compare them with the reference solution over this period. Since we work with mini-batches of training data, the summation over $B$ refers to evaluating all the cases composing the considered mini-batch. As we mentioned, we will apply (15) to the mean velocity, which is a vector field consisting of two components $\bar{\textbf{u}}=(\bar{u}^{x},\bar{u}^{y})$ in our case. Thus, we apply the previous equation to each component, resulting in the final loss as the evaluation of the following weighting loss function $\mathcal{L}$:
$\mathcal{L}(\bar{\textbf{u}}_{c},\bar{\textbf{u}}_{c,f};\theta):=w_{x}\Psi(\bar{u}^{x}_{c},\bar{u}^{x}_{c,f})+w_{y}\Psi(\bar{u}_{c}^{y},\bar{u}_{c,f}^{y}),$ (16)
where we set the weighting factors $w_{x}$ and $w_{y}$ based on each component’s loss of the baseline solver. Figure 4 outlines the algorithm and clarifies the relation between these operators.
@@
Another difference compared with the traditional approach is that we are not taking the nearest-neighboring cells when building the stencil around a face. This means that the matrix of coefficients described in Section 3.3 is no longer sparse with 5 non-zero entries per row/column, which may induce stability problems during the resolution of the system. In contrast, we follow a Deferred Correction [43] approach, commonly used for implementing high-order (HO) schemes. The technique consists of constructing the matrix of coefficients of (7) using the Upwind differencing scheme while inserting on the r.h.s. of this equation the difference between the face velocity computed through the high-order scheme and the face velocity obtained using the Upwind scheme, computed explicitly. Thus, we define the convective flux using our model as:

### v3

@@ Abstract
We propose a method for reducing the spatial discretization error of coarse computational fluid dynamics (CFD) problems by enhancing the quality of low-resolution simulations using deep learning. We feed the model with fine-grid data after projecting it to the coarse-grid discretization. We substitute the default differencing scheme for the convection term by a feed-forward neural network that interpolates velocities from cell centers to face values to produce velocities that approximate the down-sampled fine-grid data well. The deep learning framework incorporates the open-source CFD code OpenFOAM, resulting in an end-to-end differentiable model. We automatically differentiate the CFD physics using a discrete adjoint code version. We present a fast communication method between TensorFlow (Python) and OpenFOAM (c++) that accelerates the training process. We applied the model to the flow past a square cylinder problem, reducing the error from 120% to 25% in the velocity for simulations inside the training distribution compared to the traditional solver using an x8 coarser mesh. For simulations outside the training distribution, the error reduction in the velocities was about 50%. The training is affordable in terms of time and data samples since the architecture exploits the local features of the physics.
@@ 1 Introduction
Over the last decade, machine learning (ML) has arisen as a crucial technology for enhancing the solving of partial differential equations (PDEs) [1]. The field of fluid dynamics has been part of this revolution. Computational fluid dynamics (CFD) witnessed the introduction of novel learning techniques [2]. Non-linear PDEs describe fluid motion, usually showing chaotic spatiotemporal dependencies and structures with a wide range of scales, making Direct Numerical Simulation (DNS) unfeasible in most cases. More economical alternatives exist, like Large Eddy Simulations (LES) [3, 4], in which only grid structures are solved while the rest are modeled, or Reynolds-Averaged Navier-Stokes (RANS) [5], in which all the scales are modeled. However, the resolution process can be complex and time-consuming, even with LES or RANS.
@@
Enhancing CFD simulations through ML may accelerate the computations without deeply compromising accuracy or produce more accurate results with similar computational time. For this purpose, the ML model can completely substitute the traditional CFD solver, leading to reduced-order models (ROMs) or partially enhancing some specific components’ performance. The ROMs exhibit accelerated time-to-solution but pay the cost of low fidelity and a narrow range of applicability [6, 7, 8, 9]. ROMs use linear data compression techniques like dynamic-mode decomposition (DMD) [10] and principal-component analysis (PCA) [11], or non-linear compression like autoencoders [12]. When substituting a component of the traditional solver, the typical approach replaces the turbulent closure model  [13], which is one of the significant sources of error, with an ML model. Thus, Ling et al. [14] outperform the traditional solvers using a deep neural network for RANS turbulence modeling. Alternatively, Obiols-Sales et al. [15] speed up the convergence of RANS simulations by substituting the traditional solver with a subrogated model composed of a convolutional neural network (CNN) while the solution´s residual of the learned model is below a certain threshold; when this criterion is unsatisfied, they switch again to the traditional solver. Ajuria et al. [16] accelerate the algorithm by substituting the Jacobi solver with a CNN for the Poisson equation, one of the most expensive parts of the traditional numerical solver.
@@
A relevant point in ML computations is the generalization properties of the resulting approximation. Ideally, an ML model should enhance the simulation performance within multiple initial/boundary conditions, be accurate for unseen CFD simulations, and use an acceptable amount of data and time for training. Thus, Kim et al. [17] seek to improve the precision of the Finite Volume Method (FVM) [18] combining predictions of a pre-trained neural network [19] and of a traditional solver. Another promising approach consists on reducing the discretization error on coarse meshes [20], applying a method called learned discretizations [21]. For example, Kochkov et al. [22] estimate spatio-temporal operators on low-resolution grids through CNNs for accelerating DNS and LES simulations, learning from fine-resolution data. This technique calls the CFD solver during training in a solver-in-the-loop [23] ML model. This coupling demands an end-to-end differentiable framework; thus, translating the fluid solver to an automatic differentiable language is a stringent requirement [24]. This limitation reduces the applicability of this kind of method to cases employing the already implemented physics models. Another drawback of the methodology is the demanding training process from a user point of view, because CNNs require a large dataset for predicting unsteady fluid flows [25].
@@
We extend the learned discretization technique to a real case with non-trivial geometries and more complex flow dynamics, such as the flow past a square cylinder problem [26]. In these convection-dominated flows, the discretization of the convective term is a critical source of error where the differencing scheme relating the cell´s face and center values plays a major role in stability and precision. Numerous scheme versions may be found in the literature depending on the conditions of the simulated case. We reduce the spatial discretization error on coarse meshes for unsteady RANS [5] simulations by replacing the traditional differencing scheme with a feed-forward neural network, which learns from high-resolution data, which is previously restricted to the coarse grid discretization.
@@
In contrast to previous works, we embed the open-source CFD code OpenFOAM [27] in the training loop to exploit the vast range of physics and numerical solvers this program has implemented. Besides, our neural network’s architecture exploits the local features of the flow to speed up the training process. We also develop a communication method between the ML framework and the CFD code to accelerate the training process compared with a hard-disk communication. Our algorithm applies the technique to our case study, but it is easily generalisable to any fluid case simulated in OpenFOAM.
@@
Our algorithm imitates a machine learning technique commonly applied in the image processing field called super-resolution [28], following the idea of enhancing a low-quality image. Super-resolution consists of finding a mapping to reconstruct the high-resolution flow fields from the low-resolution data. Multiple methods have been proposed in the field of fluid dynamics [29] varying the architecture, the flow conditions, and the loss function for imposing the physics. In contrast, our method learns a correction while keeping the coarse-grid discretization. The method increases the accuracy of low-quality simulations online based on high-quality data, previously projected into the coarse discretization, employing data-driven models. Our model enforces physical constraints of the governing equations while enhancing certain parts of the traditional solver. Its main limitation resides in the necessity of data with an image format, that is, structured data generated from Cartesian meshes with cells of the same size.
@@ 2 Problem description
2.1 Fine discretization
The flow past a square cylinder is a well-known engineering problem with multiple industrial applications in offshore platforms [30], bridge piers [31], high-rise buildings [32], or coastal engineering [33]. The state-of-the-art of this problem is well-developed with numerous numerical [34, 35, 36] and experimental studies [37, 38, 39], leading to a case that balances industry and research. There is no general convention between authors on the definition of the computational domain and boundary conditions; thus, we define our 2D problem (see Figure 1) following [40], where $D$ is the side of the square column.
@@ 2.1 Fine discretization
According to the non-dimensional Reynolds number, which measures the ratio between inertial and viscous forces, we can identify different flow regimes [40, 34]. Most applications are interested in transition to turbulence or fully turbulent regimes, where attached vortex shedding appears in the wake of the obstacle and forms the von Kármán street vortex effect. These streets can be more or less stable depending on the turbulence grade of the flow, resulting in unsteady numerical simulations.
We fix the diameter of the square column $D=1$ m, an inlet velocity of 1 m/s, and a kinematic viscosity $\nu=2\cdot 10^{-6}$ m/s2, obtaining a Reynolds number $Re=5\cdot 10^{5}$. Although this regime corresponds to a fully turbulent flow, we will obtain results closer to the laminar wake regime due to the use of unsteady RANS equations. In the case of simulating the flow in a high-fidelity manner, it would exhibit much finer length scales [41]. Nevertheless, building a mesh that is fine enough to describe the physics with fidelity is necessary. We define our fine spatial discretization $f$ as a structured mesh formed by 303,146 quadrilateral cells of the same size. The cell size of the mesh is selected to adequately recover the von Kármán effect, although at an increased computational cost. We establish a fixed temporal step of 0.005 s, obtaining a Courant-Friedrichs-Lewy (CFL) number [42] equals 0.32. Although the simulated fine timestep is 0.005 s, we save the simulation result for each coarse timesteps (0.04 s) to force a time matching with the x8 coarse discretization. We simulate over 20 s and discard the first 4 s, obtaining 400 snapshots. We discard these steps to eliminate the simulation´s undesirable and unphysical behaviour prior to the wake formation.
@@
From now on, we will present different operators and variables with the subscripts $c$ (coarse) or $f$ (fine), indicating to which spatial discretization we apply them. For the transformation of fine data into coarse dimensions, we introduce the projection operator $\mathcal{P}_{f}^{c}:\mathbb{R}^{m_{f}}\rightarrow\mathbb{R}^{m_{c}}$, which is applied to a generic scalar-valued field $\phi$ leading to:
$\phi_{c,f}=\mathcal{P}_{f}^{c}(\phi_{f})=\frac{1}{R}\sum_{j=1}^{R}\phi_{f}^{i}(j)\qquad\text{with $i\in\{1,...,m_{c}\}$},$ (1)
@@
where the subscript $c,f$ denotes the fine data projected onto the coarse dimension, $R$ is the number of fine cells inside a coarse cell $i$, and $m_{c}$ is the number of cells of the destination coarse mesh. This operator runs through each cell of the coarse mesh and computes the arithmetic mean of the fine data contained within the considered cell. With this conservative averaging, we ensure that the resulting projected field fulfills continuity.
2.2 Coarse discretization
By down-sampling the fine mesh isotropically by a factor of x8, we define our coarse discretization $c$, which produces low-accuracy results. We apply our method to this mesh to enhance its solution from a learned model fed by highly-accurate data. We mantain the same CFL number of the fine case, that is 0.32; thus, coarse timestep is eight times bigger than the fine one. This mesh is composed by 4736 cells and 9296 interior faces. In order to illustrate the difference between both discretizations, we present Figure 2. It depicts a general comparison between the results obtained using both discretizations, the down-sampled fine result, and the error of the coarse simulation compared with the down-sampled case.
(a) Result fine discretization.
(b) Result x8 coarse discretization.
(c) Down-sampled fine result.
(d) Error (%) between (b) and (c)
Figure 2: Qualitative comparison of the results employing the different discretizations at $t=20$ s.
2.3 Generalization cases
We evaluate the generalization capabilities of our model testing its performance in two new cases derived from the one presented in Figure 1. The first case consists of extending the simulation in time out of the training distribution. Thus, we let the model evolve from 20 s to 25 s, along 125 timesteps. The second experiment consists of extending the flow past a square cylinder in domain, inserting a second column and generating the two-columns-in-tandem case. The boundary and the flow conditions are maintained from the single-column case, as well as the size of the columns. Figure 3 describes this case.
@@ 3.1 Governing equations
The system of PDEs describing the dynamics of an incompressible fluid flow is called the Navier-Stokes equations (NSE). We solve a Reynolds-averaged version of these equations, where all turbulence scales are modeled. Thus, the instantaneous values of the involved fields are decomposed into a mean component and a fluctuating one:
@@
3.2 Solver algorithm
The complex PDE system describing the problem’s physics requires an ad hoc iterative solving process. Multiple optimized algorithms exist depending on the fluid flow conditions; we use an operator-splitting method called Pressure IMplicit for Pressure-Linked Equations (PIMPLE [18]) for this case. This method is a predictor-corrector algorithm, formed by four main stages as described in Figure 4: (Stage 1) the velocity predictor, (Stage 2) the pressure equation stage, (Stage 3) the velocity corrector, and (Stage 4) the turbulence equations stage.
@@ 3.2 Solver algorithm
First, we construct a linearized momentum equation derived from (4) using the FVM and taking as input the variables at time $t$. The solution of this equation is the predicted mean velocity $\bar{\textbf{u}}^{*}$, which does not satisfy continuity (3). During the second step, we form and solve a Poisson equation by combining (3)-(4), leading to a predicted mean pressure $\bar{p}^{*}$. During (Stage 3), the algorithm corrects the mean velocity based on the expected pressure computed in the previous step. This stage is repeated in the correction loop until the corrected mean velocity $\bar{\textbf{u}}^{**}$ is conservative. At the end of this stage, we obtain a candidate mean velocity and a candidate mean pressure, which can be sent back to (Stage 1) in case the convergence criteria are not satisfied. In that case, we would be restarting the PIMPLE loop taking $\bar{\textbf{u}}^{**}$ and $\bar{p}^{*}$ as the new input of the iteration. Once convergence is reached, we obtain the mean velocity and mean pressure for the time $t+1$. Finally, we jump into the final stage, where we solve the turbulence equations (5)-(6), producing $k$ and $\omega$ at time $t+1$.
@@ 3.3 Differencing schemes
The first step of the solver consists of linearizing the momentum equation to obtain a system of linear equations, which can be generalized for a cell $p$ as:
@@
where $N$ is the selected stencil around face $F$ and $w$ are the interpolation weights. The mean face velocity is not computed explicitly; in contrast, the interpolation weights $w_{i}$ contribute to the matrix of coefficients through $a_{p}$ and $a_{nb}$ of (7), following an implicit approach. A differencing scheme is an algebraic expression for relating a cell’s face and center values based on some constant interpolation weights. Experts have developed numerous differencing schemes (e.g., Upwind, QUICK, TVD schemes [18]) depending on the flow condition, resulting in variations of the stencil size and the interpolation weights. However, no general approach has been reached. For this reason, its choice is case-dependent, producing different results in accuracy and stability. Our approach builds a custom version of a differencing scheme employing deep learning.
@@ 4.1 Formal description
As mentioned in Section 2, our problem is fully turbulent, so the convective contribution is the most crucial term of (4). The use of an appropriate mesh is essential for discretizing this term since a coarser mesh would considerably impact the accuracy due to its spatial discretization error. In particular, when we discretize the convective term using FVM, we obtain the following expression:
@@
As explained, the critical interpolation is carried out by a differencing scheme. Our approach substitutes the differencing scheme used for the convective term with a custom version based on data-driven discretizations [21]. Thus, we replace the constant interpolation weights $w_{i}$ of (8) by a vector of optimized interpolation weights $\widetilde{\textbf{w}}_{i}$ generated by a feed-forward neural network. We can therefore define a modified inverse operator ~$\widetilde{\mathcal{I}}:\mathbb{R}^{3N}\rightarrow\mathbb{R}^{2}$, such that:
@@
where $\theta$ are the parameters of the neural network. In contrast to traditional differencing schemes, our approach takes the mean velocity and mean pressure as inputs and returns a vector of interpolation weights that are space and time-dependent. The resulting operators $\mathcal{I}$ and $\widetilde{\mathcal{I}}$ are local, meaning they compute the value at face $F$ based only on its neighborhood. However, we apply the local operators to all the internal faces of the domain, resulting in the mean face velocities of the domain $\bar{\textbf{u}}_{F(n)}$ as:
@@
$\textbf{s}^{t}:=\bigcup^{m}_{j=1}\{\bar{\textbf{u}}^{t}_{(j)},\bar{p}^{t}_{(j)},k^{t}_{(j)},\omega^{t}_{(j)}\},$ (13)
being $m$ the number of cells of the mesh. We fully define our model by composing these operators ($\mathcal{F}\circ\mathcal{I})(\textbf{s}^{t})$ and applying them to a spatial discretization. In particular, we run the learned model on coarse meshes, which we indicate by the subscript $c$. We train it in such a way that the generated interpolation weights deliver a mean velocity $\bar{\textbf{u}}_{c}$ as close as possible to the mean velocity from the fine discretization projected onto the coarse one $\bar{\textbf{u}}_{c,f}$. Thus, we can formulate the training as the following minimization problem:
@@
Our method, called Deep-Convection (DC), mimics the super-resolution techniques employed in image processing, which augment the accuracy of low-quality images based on high-quality data [28]. However, in our approach the enhancement is carried out maintaining the resolution and previously projecting the training data to the coarse-grid discretization. With this approach, the primary benefit is that it maintains the physical constraints of the governing equations while exploiting the benefits of a data-driven model for enriching certain parts of the traditional solver with high-resolution data.
@@
where $\phi_{c}$ is a general scalar-valued field computed using the coarse discretization and $\phi_{c,f}$ is the projection of the fine data $\phi_{f}$ onto the coarse mesh $c$. In this case, $\phi_{c,f}$ is the training data or the ground truth. Due to its multi-step feature, we gather the solver prediction results over a number of accumulated timesteps $T$ and then compare them with the reference solution over this period. The number of accumulated timesteps $T$ is a crucial hyperparameter, which should be selected based on a trial-and-error procedure. A higher T will imply a better generalization of the model, but it will increase the training process cost. Different features of the problem, such as the complexity of the fluid flow, the number of interior faces of the domain, or the number of timesteps outside of the training distribution you pretend to simulate, will influence the value of T, so it is complicated to give an a priori estimation. Since we work with mini-batches of training data, the summation over $B$ refers to evaluating all the cases composing the considered mini-batch. As we mentioned, we apply (15) to the mean velocity, which is a vector field consisting of two components $\bar{\textbf{u}}=(\bar{u}^{x},\bar{u}^{y})$ in our case. Thus, we employ the previous equation with each component, resulting in the final loss as the evaluation of the following weighting loss function $\mathcal{L}$:
$\mathcal{L}(\bar{\textbf{u}}_{c},\bar{\textbf{u}}_{c,f};\theta):=\kappa_{x}\Psi(\bar{u}^{x}_{c},\bar{u}^{x}_{c,f})+\kappa_{y}\Psi(\bar{u}_{c}^{y},\bar{u}_{c,f}^{y}),$ (16)
where we set the weighting factors $\kappa_{x}$ and $\kappa_{y}$ based on each component’s loss of the baseline coarse solver. We construct the loss function based just on the velocities for the sake of simplicity and with the intention of obtaining an indirect improvement on the rest of the variables [22]. Including more fields in the loss function increases the training cost due to the reconstruction of additional intermediate derivatives, and generates a multi-objective optimization problem of coupled variables, commonly difficult to address [46]. Another possibility could have been to include the correction of other variables, adding them to the neural network’s output while maintaining the loss function. This solution would also induce a more costly training process associated with a more complex optimization problem. A solution could be to introduce independent extra neural networks related to the correction of the rest of the variables and train them separately in a sequential way. However, exploring these alternatives is beyond the scope of the present paper, we may consider them in future works. Figure 5 outlines the algorithm and clarifies the relation between these operators. Thus, we start with the input variables at time $t$, then we apply sequentially the operators $\widetilde{\mathcal{I}_{c}}$ and $\mathcal{F}_{c}$, and finally obtain as output all the variables at time $t+1$. Then we apply $\mathcal{L}$ to $\bar{\textbf{u}}_{c}^{t+1}$ to get the corresponding loss.
@@
Another difference compared with the traditional approach is that we are not taking the nearest-neighboring cells when building the stencil around a face. This means that the matrix of coefficients described in Section 3.3 is no longer sparse with 5 non-zero entries per row/column, which may induce stability problems during the iterative procedure for solving the system. In contrast, we follow a Deferred Correction [47] approach, commonly used for implementing high-order (HO) schemes. The technique consists of constructing the matrix of coefficients of (7) using the Upwind differencing scheme while inserting on the r.h.s. of this equation the difference between the face velocity computed through the high-order scheme and the face velocity obtained using the Upwind scheme, computed explicitly. Thus, we define the convective flux using our model as: